Find two consecutive intergers such that the sum of 2 times the first interger and 7 times second interger is 88.
Answer by checkley75(3666)

LET X BE THE FIRST INTERGER & X=1 BE THE SECOND.
2X+7(X+1)=88
2X+7X+7=88
9X=88-7
9X=81
X=81/9
X=9 ANSWER FOR THE FIRST INTERGER.
X+1=10 FOR THE SECOND INTERGER.
PROOD
2*9+7*10=88
18+70=88
88=88